OpenStudy (anonymous):

solve the system of equations using any method? x+y=7 3x-2y=11

OpenStudy (anonymous):

@spotless ??

OpenStudy (anonymous):

@samalami ??

OpenStudy (anonymous):

@sami-21 ??

OpenStudy (anonymous):

@shafqat_uet helpp??

OpenStudy (anonymous):

no one no's how to do this??

OpenStudy (anonymous):

first solve equation 1 for y we get x+y=7 y=7-x put this in the second equation 3x-2(7-x)=11 3x-14+2x=11 5x=11+14 5x=25 x=5 put this in any of the above given equations you will get y value.

OpenStudy (anonymous):

\[y =2\]

OpenStudy (anonymous):

show work @shafqat_uet

OpenStudy (anonymous):

x=5 and y=2

OpenStudy (anonymous):

@shafqat_uet i asked him to do this.

OpenStudy (anonymous):

so it would be 5+y=7?? @sami-21

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

solve this now you will get y

OpenStudy (anonymous):

how do you do this??

OpenStudy (anonymous):

5x=25 after distributed so x=5

OpenStudy (anonymous):

and y=2

OpenStudy (anonymous):

@broderick365 we have 5+y=7 now subtract 5 from both sides 5-5+y=7-5 y=2

OpenStudy (anonymous):

thanks!!!